On January 1, Year 1, Jing Company purchased office equipment that cost $16,200 cash. The equipment was delivered under terms FOB shipping point, and transportation cost was $2,300. The equipment had a five-year useful life and a $6,040 expected salvage value.

Assuming the company uses the double-declining-balance depreciation method, what are the amounts of depreciation expense and accumulated depreciation, respectively, that would be reported in the financial statements prepared as of December 31, Year 3?
Answer #1

Cost of equipment = 16200+2300 = 18500

Dep rate = 100/5*2 = 40%

Year 1 Dep = 18500*40% = 7400

Year 2 Dep = 18500*60%*40% = 4440

Year 3 Dep = (18500-7400-4440-6040) = 620

Year 3 Depreciation expense = $620

Accumulated dep = 7400+4440+620 = $12460